# Contraherent cosheaves of contramodules on Noetherian formal schemes

### 29 Mar 2026 v1

@@ 2.7. Tensor products of adic topological rings
Put $W=T\otimes_{R}S$. Let $J\subset S$ be an ideal of definition and $K\subset T$ be an ideal of definition. Since the intersection of any two ideals of definition in $R$ is also an ideal of definition in $R$,  Lemma 2.1.3(3) implies existence of an ideal of definition $I\subset R$ such that $f(I)\subset J$ and $g(I)\subset K$. Then, according to Lemma 2.7.1(d),  $L=W(f^{\prime}(K)+g^{\prime}(J))\subset W$ is an ideal of definition in $W$. We have $W/L\simeq T/K\otimes_{R/I}S/J$.
@@ 2.8. Formal open immersions and formal open coverings
Let us say that a homomorphism of (discrete) commutative rings $f\colon R\longrightarrow S$ is an open immersion if the induced map of the spectra $\operatorname{Spec}S\longrightarrow\operatorname{Spec}R$ is an open immersion of affine schemes. Equivalently, $f$ is an open immersion if and only if it is a flat ring epimorphism of finite presentation [39, Section XI.2], [10, Theoreme IV.17.9.1]. In particular, we will say that a homomorphism of commutative rings $R\longrightarrow S$ is a principal open immersion if there exists an element $r\in R$ such that $S$ is isomorphic to $R[r^{-1}]$ as a commutative $R$-algebra.
@@
We are not aware of any argument within the realm of commutative ring theory. The following proof is, basically, a collection of references. By [10, Theorem IV.18.1.2], the functor $X\longmapsto\operatorname{Spec}R/I\times_{\operatorname{Spec}R}X$ is an equivalence between the category of étale schemes over $\operatorname{Spec}R$ and the category of étale schemes over $\operatorname{Spec}R/I$. By [10, Theorem IV.17.9.1], an étale morphism is an open immersion if and only if it is radicial (in the sense of [11, Section I.3.7]). It follows easily that an étale morphism $X\longrightarrow\operatorname{Spec}R$ is an open immersion if and only if the corresponding morphism $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}X\longrightarrow\operatorname{Spec}R/I$ is an open immersion. So the map $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}{-}$ is an equivalence between the category of open subschemes in $\operatorname{Spec}R$ and the category of open subschemes in $\operatorname{Spec}R/I$. The same conclusion can be simply obtained from the facts that open subschemes in schemes correspond bijectively to open subsets in the underlying topology, and the underlying topological spaces of $\operatorname{Spec}R$ and $\operatorname{Spec}R/I$ coincide.
The key step is to prove that an open subscheme $Y\subset\operatorname{Spec}R$ is affine if and only if the scheme $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$ is affine. The “only if” assertion is obvious. To prove the “if”, notice that the underlying topological spaces of $Y$ and $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$ coincide. Quasi-compactness and quasi-separatedness are properties of the underlying topological space of a scheme; and all affine schemes are quasi-compact and quasi-separated. So the scheme $Y$ is quasi-compact and quasi-separated (in fact, separated, as an open subscheme of an affine scheme). By [8, Théorème II.5.2.1(d)], [9, IV.1.7.17], [13, Lemmas Tags 01XB and 01XF or 01XG], a quasi-compact scheme $Y$ is affine if and only if $H^{1}(Y,\mathcal{F})=0$ for every quasi-coherent sheaf $\mathcal{F}$ on $Y$. Here $H^{*}(Y,{-})$ denotes the cohomology groups of sheaves of abelian groups on $Y$, which only depend on the underlying topological space of $Y$. It remains to point out that every quasi-coherent sheaf $\mathcal{F}$ on $Y$ has a finite decreasing filtration by quasi-coherent subsheaves $\mathcal{F}\supset I\mathcal{F}\supset I^{2}\mathcal{F}\supset\dotsb\supset I^{n}\mathcal{F}=0$, where $n\geq 1$ is an integer such that $I^{n}=0$ in $R$, and the successive quotient sheaves $I^{m}\mathcal{F}/I^{m+1}\mathcal{F}$ are quasi-coherent sheaves on $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$.
@@
We will say that a collection of tight continuous maps of adic topological rings $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ is a formal open covering if it satisfies the equivalent conditions of Lemma 2.8.10. Clearly, any formal open covering of an adic topological ring has a finite subcovering. A principal formal open covering is a formal open covering $(f_{\alpha})$ such that every map $f_{\alpha}$ is a principal formal open immersion.

### 17 Apr 2026 v2

@@ 2.7. Tensor products of adic topological rings
Put $W=T\otimes_{R}S$. Let $J\subset S$ be an ideal of definition and $K\subset T$ be an ideal of definition. Since the product of any two ideals of definition in $R$ is also an ideal of definition in $R$,  Lemma 2.1.3(3) implies existence of an ideal of definition $I\subset R$ such that $f(I)\subset J$ and $g(I)\subset K$. Then, according to Lemma 2.7.1(d),  $L=W(f^{\prime}(K)+g^{\prime}(J))\subset W$ is an ideal of definition in $W$. We have $W/L\simeq T/K\otimes_{R/I}S/J$.
@@ 2.8. Formal open immersions and formal open coverings
Let us say that a homomorphism of (discrete) commutative rings $f\colon R\longrightarrow S$ is an open immersion if the induced map of the spectra $\operatorname{Spec}S\longrightarrow\operatorname{Spec}R$ is an open immersion of affine schemes. Equivalently, $f$ is an open immersion if and only if it is a flat ring epimorphism of finite presentation [40, Section XI.2], [11, Theoreme IV.17.9.1]. In particular, we will say that a homomorphism of commutative rings $R\longrightarrow S$ is a principal open immersion if there exists an element $r\in R$ such that $S$ is isomorphic to $R[r^{-1}]$ as a commutative $R$-algebra.
@@
We are not aware of any argument within the realm of commutative ring theory. The following proof is, basically, a collection of references. By [11, Theorem IV.18.1.2], the functor $X\longmapsto\operatorname{Spec}R/I\times_{\operatorname{Spec}R}X$ is an equivalence between the category of étale schemes over $\operatorname{Spec}R$ and the category of étale schemes over $\operatorname{Spec}R/I$. By [11, Theorem IV.17.9.1], an étale morphism is an open immersion if and only if it is radicial (in the sense of [12, Section I.3.7]). It follows easily that an étale morphism $X\longrightarrow\operatorname{Spec}R$ is an open immersion if and only if the corresponding morphism $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}X\longrightarrow\operatorname{Spec}R/I$ is an open immersion. So the map $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}{-}$ is an equivalence between the category of open subschemes in $\operatorname{Spec}R$ and the category of open subschemes in $\operatorname{Spec}R/I$. The same conclusion can be simply obtained from the facts that open subschemes in schemes correspond bijectively to open subsets in the underlying topology, and the underlying topological spaces of $\operatorname{Spec}R$ and $\operatorname{Spec}R/I$ coincide.
The key step is to prove that an open subscheme $Y\subset\operatorname{Spec}R$ is affine if and only if the scheme $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$ is affine. The “only if” assertion is obvious. To prove the “if”, notice that the underlying topological spaces of $Y$ and $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$ coincide. Quasi-compactness and quasi-separatedness are properties of the underlying topological space of a scheme; and all affine schemes are quasi-compact and quasi-separated. So the scheme $Y$ is quasi-compact and quasi-separated (in fact, separated, as an open subscheme of an affine scheme). By [9, Théorème II.5.2.1(d)], [10, IV.1.7.17], [14, Lemmas Tags 01XB and 01XF or 01XG], a quasi-compact scheme $Y$ is affine if and only if $H^{1}(Y,\mathcal{F})=0$ for every quasi-coherent sheaf $\mathcal{F}$ on $Y$. Here $H^{*}(Y,{-})$ denotes the cohomology groups of sheaves of abelian groups on $Y$, which only depend on the underlying topological space of $Y$. It remains to point out that every quasi-coherent sheaf $\mathcal{F}$ on $Y$ has a finite decreasing filtration by quasi-coherent subsheaves $\mathcal{F}\supset I\mathcal{F}\supset I^{2}\mathcal{F}\supset\dotsb\supset I^{n}\mathcal{F}=0$, where $n\geq 1$ is an integer such that $I^{n}=0$ in $R$, and the successive quotient sheaves $I^{m}\mathcal{F}/I^{m+1}\mathcal{F}$ are quasi-coherent sheaves on $\operatorname{Spec}R/I\times_{\operatorname{Spec}R}Y$.
@@
We will say that a collection of tight continuous maps of adic topological rings $f_{\alpha}\colon R\longrightarrow S_{\alpha}$ is a formal open covering if it satisfies the equivalent conditions of Lemma 2.8.11. It is clear from Lemma 2.8.11(3) that any formal open covering of an adic topological ring has a finite subcovering. A principal formal open covering is a formal open covering $(f_{\alpha})$ such that every map $f_{\alpha}$ is a principal formal open immersion.
@@
Let $R$ and $S$ be adic topological rings, let $f\colon R\longrightarrow S$ be a continuous ring map, and let $g_{\alpha}\colon R\longrightarrow T_{\alpha}$ be a formal open covering. Put $W_{\alpha}=T_{\alpha}\otimes_{R}S$, and denote by $f^{\prime}_{\alpha}\colon T_{\alpha}\longrightarrow W_{\alpha}$ and $g^{\prime}_{\alpha}\colon S\longrightarrow W_{\alpha}$ the induced ring maps. Endow the commutative rings $W_{\alpha}$ with the tensor product topologies. Then the collection of ring maps $g^{\prime}_{\alpha}\colon S\longrightarrow W_{\alpha}$ is a formal open covering, too.
Proof.
The argument, based on the fact that any base change of an open covering (of discrete commutative rings) is an open covering, is straightforward. ∎
Lemma 2.8.14.
@@
2.9. Locality of Noetherianity
The following lemma is well-known.
@@ 2.9. Locality of Noetherianity
(a) Let $f\colon R\longrightarrow S$ be an open immersion of commutative rings. Assume that the ring $R$ is Noetherian. Then the ring $S$ is Noetherian.
(b) Let $f\colon R\longrightarrow S_{\alpha}$ be an open covering of a commutative ring $R$. Assume that all the rings $S_{\alpha}$ are Noetherian. Then the ring $R$ is Noetherian.